# Fully viable DHOST bounce with extra scalar

### v1

@@ 1 Introduction
The aim of the this paper is to build a novel explicit bouncing scenario within the exceptional subclass of DHOST Ia theories such that
@@ 2.1 DHOST Ia with extra scalar
Here $\phi_{,\mu}\equiv\partial_{\mu}\phi$, $X\equiv g^{\mu\nu}\phi_{,\mu}\phi_{,\nu}$, $\chi_{,\mu}\equiv\partial_{\mu}\chi$, $Y\equiv g^{\mu\nu}\chi_{,\mu}\chi_{,\nu}$, and $L_{I}$ ($I=1,2,3,4,5$) are
@@
where the subscript $X$ denotes the derivative with respect to $X$. We emphasize that the coupling of $\chi$ to the DHOST theory does not spoil the degeneracy, as the scalar field $\chi$ does not feature higher derivatives and the coupling terms do not involve derivatives of $\phi$ Deffayet:2020ypa ; Mironov:2024xxx . Hence, the whole system does not suffer from the Ostrogradsky ghost.
@@ 2.2 Quadratic action for perturbations
In order to deal with the instability and superluminality issue we go to the perturbation level. The perturbed spatially-flat Friedmann-Lemaître-Robertson-Walker (FLRW) metric is given by
$ds^{2}=(1+\alpha)^{2}dt^{2}-a^{2}e^{2\zeta}\left(\delta_{ij}+h_{ij}^{T}+\frac{1}{2}h_{ik}^{T}h_{j}^{k\;T}\right)(dx^{i}+\partial^{i}\beta\;dt)(dx^{j}+\partial^{j}\beta\;dt),$ (4)
where $\alpha,\beta$ and $\zeta$ denote lapse, shift and curvature perturbations, respectively, and $h_{ij}^{T}$ is traceless and transverse perturbation. Throughout this paper we work in the unitary gauge $\delta\phi=0$. The lapse and shift perturbations $\alpha$ and $\beta$ appear in the quadratic action as Lagrange multipliers, so they can be integrated out. After all, we obtain a quadratic action $S^{(2)}=S^{(2)}_{\text{tensor}}+S^{(2)}_{\text{scalar}}$ for the perturbations $h_{ij}^{T}$, $\zeta$ and $\delta\chi$, where
@@
$f=$ $\;-\frac{\mathcal{G}_{\mathcal{T}}+\mathcal{D}\dot{\phi}+\mathcal{F}_{\mathcal{T}}\Delta}{\Theta},$ (10b)
$g=$ $\;-\frac{\mathcal{G}_{\mathcal{T}}}{\Theta}\left(1-3\frac{P_{Y}}{Q}\Delta\right),$ (10c)
$\mathcal{G}_{\mathcal{S}}=$ $\;\frac{\Sigma\mathcal{G}_{\mathcal{T}}^{2}}{\Theta^{2}}+3\mathcal{G}_{\mathcal{T}},$ (10d)
$\mathcal{F}_{\mathcal{S}}=$ $\;\frac{1}{a}\frac{d}{dt}\left(\frac{a\mathcal{G}_{\mathcal{T}}(\mathcal{G}_{\mathcal{T}}+\mathcal{D}\dot{\phi}+\mathcal{F}_{\mathcal{T}}\Delta)}{\Theta}\right)-\mathcal{F}_{\mathcal{T}},$ (10e)
$\mathcal{D}=$  $\;2\dot{\phi}(F_{2X}-A_{1}),$ (10f)
$\Delta=$ $\;\frac{X}{2\mathcal{G}_{\mathcal{T}}}(2A_{1}-4F_{2X}-A_{3}X),$ (10g)
@@
$+\frac{1}{2}H\dot{\phi}^{4}\left(8A_{1X}+3A_{3}\right)+\dot{\phi}^{5}\ddot{\phi}\left(A_{5}-A_{3X}\right),$ (10h)
@@
$+\dot{\phi}^{7}\left(-3H\left(A_{3\phi X}+2A_{5X}\ddot{\phi}\right)-2A_{5X}\dddot{\phi}\right)-2\dot{\phi}^{8}\ddot{\phi}\left(A_{5XX}\ddot{\phi}+A_{5\phi X}\right).$ (10i)
@@ 2.3 Stabilities and non-superluminality conditions: an exceptional subclass
The stability and non-superluminality conditions for the tensor sector immediately follow from the action (5) that
$\mathcal{G}_{\mathcal{T}}\geq\mathcal{F}_{\mathcal{T}}>0,$ (11)
while those for the scalar sector require some involved algebra. First, the stability conditions amount to the positive definiteness of kinetic matrices (9c), i.e.
@@ 3 DHOST Sector
In this section we consider the decoupled case (i.e., $\dot{\chi}=0$) to construct a fully healthy model in the exceptional subclass of DHOST Ia, in which the independent Lagrangian functions are $F(\phi,X)$, $F_{2}(\phi,X)$, $K(\phi,X)$ and $A_{1}(\phi,X)$. To this end, we adopt the “inverse method” used in Ijjas:2016tpn ; Ijjas:2016vtq ; Libanov:2016kfc ; Kolevatov:2017voe ; Cai:2017dyi ; Mironov:2018oec to choose a specific form of the Hubble parameter $H(t)$ and reconstruct the Lagrangian functions $F(\phi(t),X(t))$, $F_{2}(\phi(t),X(t))$, $K(\phi(t),X(t))$ and $A_{1}(\phi(t),X(t))$ that obey the background equations of motion and satisfy the constraint (22).
@@ 3.2 Stability and non-superluminality
where the function $\bar{\Sigma}$ given in (10i) is reduced to
@@ 4.3 Conversion of entropic perturbations into curvature fluctuations
$b=0.0158,\quad\bar{t}_{c}=10\tau=100,\quad q=\frac{1}{12},\quad\lambda=0.35,\quad\bar{\chi}_{0}=\frac{1}{10}.$ (77b)
Fig. 10 shows the result of the numeric computation, where the initial time is $\bar{t}_{i}=-3\times 10^{3}$. The conversion process takes place around $4\tau=\bar{t}_{c}-6\tau\lesssim\bar{t}\lesssim\bar{t}_{c}+6\tau=16\tau$, during which about $98\%$ of $\delta\bar{\chi}_{k}$ is converted into $\tilde{\zeta}_{k}$ and the comoving Hubble radius $(aH)^{-1}$ decreases by order of one e-fold. This implies that the conversion is smooth and highly efficient, cf. Fertig:2016czu .

### v2

@@ 1 Introduction
The aim of this paper is to build a novel explicit bouncing scenario within the exceptional subclass of DHOST Ia theories such that
@@ 2.1 DHOST Ia with extra scalar
Here $X\equiv g^{\mu\nu}\phi_{,\mu}\phi_{,\nu}$, $\phi_{,\mu}\equiv\partial_{\mu}\phi$, $Y\equiv g^{\mu\nu}\chi_{,\mu}\chi_{,\nu}$, $\chi_{,\mu}\equiv\partial_{\mu}\chi$, and $L_{I}$ ($I=1,2,3,4,5$) are
@@
where the subscript $X$ denotes the derivative with respect to $X$. The independent functions of the model (1) are then $F(\phi,X)$, $K(\phi,X)$, $F_{2}(\phi,X)$, $A_{1}(\phi,X)$, $A_{3}(\phi,X)$ and $P(\chi,Y,\phi)$. We emphasize that the coupling of $\chi$ to the DHOST theory does not spoil the degeneracy, as the scalar field $\chi$ does not feature higher derivatives and the coupling terms do not involve derivatives of $\phi$ Deffayet:2020ypa ; Mironov:2024xxx . Hence, the whole system does not suffer from the Ostrogradsky ghost.
@@ 2.2 Quadratic action for perturbations
As mentioned in the Introduction, we aim to construct a bouncing scenario fully viable in the sense that (i) it is free of any pathologies such as ghost instability, gradient instability and superluminality and (ii) it predicts power spectra of primordial perturbations compatible with observations. On a spatially homogeneous background, the aforementioned pathologies, if present, manifest themselves in the linear perturbation theory, see e.g. Rubakov:2014jja . Hence, we consider the linear theory of the model (1).
The perturbed spatially-flat Friedmann-Lemaître-Robertson-Walker (FLRW) metric is given by
$ds^{2}=(1+\alpha)^{2}dt^{2}-a^{2}(t)^{2\zeta}\left(\delta_{ij}+h_{ij}^{TT}+\frac{1}{2}h_{ik}^{TT}h_{j}^{k\;TT}\right)(dx^{i}+\partial^{i}\beta\;dt)(dx^{j}+\partial^{j}\beta\;dt),$ (4)
where $a(t)$ is the scale factor, $\alpha,\beta$ and $\zeta$ denote lapse, shift and curvature perturbations, respectively, and $h_{ij}^{TT}$ is traceless and transverse tensor perturbation. We did not include the vector modes in the perturbed metric (4), as they are non-dynamical in scalar-tensor theories Mironov:2024xxx . [^2] Even though non-dynamical, the vector perturbations can be problematic in bouncing cosmologies, since they grow as $1/a^{2}$ during the contracting phase Battefeld:2014uga . The ekpyrotic contracting phase, which is employed in this work, smooths and flattens the universe, since the energy density of the homogeneous ekpyrotic scalar field grows faster than $1/a^{6}$, so that the vector perturbations remain negligible compared to the background. The regrowth of vector modes during the bounce phase discussed in Xue:2011nw does not occur in this work, because the overall change in the scale factor during the bounce phase is negligible, see (28) and Fig. 1. Throughout this paper we work in the unitary gauge $\delta\phi=0$. The lapse and shift perturbations $\alpha$ and $\beta$ appear in the quadratic action as Lagrange multipliers, so they can be integrated out. After all, we obtain a quadratic action $S^{(2)}=S^{(2)}_{\text{tensor}}+S^{(2)}_{\text{scalar}}$ for the perturbations $h_{ij}^{TT}$, $\zeta$ and $\delta\chi$, where
@@
$Q=$ $\;P_{Y}+2YP_{YY},$ (10b)
$f=$ $\;-\frac{\mathcal{G}_{\mathcal{T}}+\mathcal{D}\dot{\phi}+\mathcal{F}_{\mathcal{T}}\Delta}{\Theta},$ (10c)
$g=$ $\;-\frac{\mathcal{G}_{\mathcal{T}}}{\Theta}\left(1-3\frac{P_{Y}}{Q}\Delta\right),$ (10d)
$\mathcal{G}_{\mathcal{S}}=$ $\;\frac{\Sigma\mathcal{G}_{\mathcal{T}}^{2}}{\Theta^{2}}+3\mathcal{G}_{\mathcal{T}},$ (10e)
$\mathcal{F}_{\mathcal{S}}=$ $\;\frac{1}{a}\frac{d}{dt}\left(\frac{a\mathcal{G}_{\mathcal{T}}(\mathcal{G}_{\mathcal{T}}+\mathcal{D}\dot{\phi}+\mathcal{F}_{\mathcal{T}}\Delta)}{\Theta}\right)-\mathcal{F}_{\mathcal{T}},$ (10f)
$\mathcal{D}=$  $\;2\dot{\phi}(F_{2X}-A_{1}),$ (10g)
$\Delta=$ $\;\frac{X}{2\mathcal{G}_{\mathcal{T}}}(2A_{1}-4F_{2X}-A_{3}X),$ (10h)
@@
$+\frac{1}{2}H\dot{\phi}^{4}\left(8A_{1X}+3A_{3}\right)+\dot{\phi}^{5}\ddot{\phi}\left(A_{5}-A_{3X}\right),$ (10i)
@@
$+\dot{\phi}^{7}\left(-3H\left(A_{3\phi X}+2A_{5X}\ddot{\phi}\right)-2A_{5X}\dddot{\phi}\right)-2\dot{\phi}^{8}\ddot{\phi}\left(A_{5XX}\ddot{\phi}+A_{5\phi X}\right).$ (10j)
@@ 2.3 Stabilities and non-superluminality conditions: an exceptional subclass
The main types of instabilities that can potentially emerge at the linearized level of perturbations in the high-momentum regime are ghost and gradient instabilities (see Rubakov:2014jja for details). The presence of a ghost instability implies that the vacuum is quantum mechanically unstable, while the gradient instability leads to the exponential growth of perturbations.
For the tensor modes, the ghost instability appears when $\mathcal{G}_{\mathcal{T}}<0$ and $\mathcal{F}_{\mathcal{T}}<0$ in (5), while the gradient instability shows up when $\mathcal{G}_{\mathcal{T}}>0$, $\mathcal{F}_{\mathcal{T}}<0$ or $\mathcal{G}_{\mathcal{T}}<0$, $\mathcal{F}_{\mathcal{T}}>0$. The stable situation thus corresponds to $\mathcal{G}_{\mathcal{T}}>0$ and $\mathcal{F}_{\mathcal{T}}>0$. On the other hand, the non-superluminality condition, which requires the propagation speed of tensor perturbations not greater than unity, is $c_{T}^{2}\equiv\mathcal{F}_{\mathcal{T}}/\mathcal{G}_{\mathcal{T}}\leq 1$. Thus, the stability and non-superluminality conditions for the tensor sector are given by
$\mathcal{G}_{\mathcal{T}}\geq\mathcal{F}_{\mathcal{T}}>0.$ (11)
Similar arguments apply to the scalar sector, but two-field cases require some involved algebra. First, the stability conditions amount to the positive definiteness of kinetic matrices (9c) (see Mironov:2020pqh ; Mironov:2024xxx ), i.e.
@@ 3 DHOST Sector
In this section we consider the decoupled case (i.e., $\dot{\chi}=0$) to construct a fully healthy model in the exceptional subclass of DHOST Ia, in which the independent Lagrangian functions are $F(\phi,X)$, $F_{2}(\phi,X)$, $K(\phi,X)$ and $A_{1}(\phi,X)$. We remind that the Lagrangian function $A_{3}(\phi,X)$ is no longer independent in the exceptional subclass due to the condition (17). To this end, we adopt the inverse method used in Ijjas:2016tpn ; Ijjas:2016vtq ; Libanov:2016kfc ; Kolevatov:2017voe ; Cai:2017dyi ; Mironov:2018oec to choose a specific form of the Hubble parameter $H(t)$ and reconstruct the Lagrangian functions $F(\phi(t),X(t))$, $F_{2}(\phi(t),X(t))$, $K(\phi(t),X(t))$ and $A_{1}(\phi(t),X(t))$ that obey the background equations of motion and satisfy the constraint (22).
@@ 3.2 Stability and non-superluminality
where the function $\bar{\Sigma}$ given in (10j) is reduced to
@@ 4.3 Conversion of entropic perturbations into curvature fluctuations
$b=0.0158,\quad\bar{t}_{c}=10\tau=100,\quad q=\frac{1}{12},\quad\lambda=0.35,\quad\bar{\chi}_{0}=\frac{1}{10},$ (77b)
where the values of $\tau$, $u$, and $w$ are the same as those given in Mironov:2018oec , while the values of parameters $\bar{t}_{c}$, $q$, $\lambda$ and $\bar{\chi}_{0}$ are chosen based on an educated guess obtained by conducting some tentative numerical analysis. Fig. 10 shows the result of the numeric computation, where the initial time is $\bar{t}_{i}=-3\times 10^{3}$. The conversion process takes place around $4\tau=\bar{t}_{c}-6\tau\lesssim\bar{t}\lesssim\bar{t}_{c}+6\tau=16\tau$, during which about $98\%$ of $\delta\bar{\chi}_{k}$ is converted into $\tilde{\zeta}_{k}$ and the comoving Hubble radius $(aH)^{-1}$ decreases by order of one e-fold. This implies that the conversion is smooth and highly efficient, cf. Fertig:2016czu .